OpenStudy (accidentalaichan):

Which factor do you cancel from the numerator and denominator to simplify x^2 + 3x -4 / x^2 -16?

OpenStudy (solomonzelman):

\(\Large\color{red}{ x^2 + 3x -4 =(x-1)(x+4) }\) \(\Large\color{blue}{ x^2 -16 =(x-4)(x+4) }\)

OpenStudy (accidentalaichan):

x - 3

OpenStudy (accidentalaichan):

That's what I guess. x + 3 and x - 3 cancel out, so it makes sense

OpenStudy (ʕ•ᴥ•ʔ):

x+3 and x-3 don't cancel..

OpenStudy (ipwnbunnies):

(x-3)? o.o When Solomon factored both the numerator and denominator, he factored each into two binomials. (x+4) is the common factor in both the numerator and denominator. That factor will cancel, and you will be left with (x-1) in the numerator and (x-4) in the denominator.

OpenStudy (accidentalaichan):

Okay x.x Thanks, Bunni ♥